# M2CVD: Multi-Model Collaboration for Code Vulnerability Detection

### v1

@@ 5.2. RQ2: Effects of vulnerability description refinement for detection performance
b) UniXcoder: Also adopts the optimal configuration as documented in existing literature.
@@
d) M2CVD(UniXcoder) w/o PII: It omits the M2CVD comparison process. The LLM’ first vulnerability detection, when combined with the code, employs the UniXcoder model for prediction.
@@
f) M2CVD(UniXcoder): Also use the standard M2CVD process, with ChatGPT as the LLMs and UniXcoder as the code model.
The results from the experiment as shown in the table 3, which presents a comparison results of model accuracy under different M2CVD configurations. The experimental result evidence suggests a improvement in model accuracy when integrating Phase II feedback into the M2CVD framework. Notably, the M2CVD (CodeBERT) w/o PII outperforms the CodeBERT model by a margin of 1.91%. Similarly, the M2CVD (UniXcoder) w/o PII configuration outperforms the UniXcoder by a margin of 2.23%. The enhancements are more pronounced when the comparison includes Phase II, as observed with the M2CVD (CodeBERT) configuration, which incorporates Phase II feedback, outperforms the CodeBERT model by a margin of 2.51%. And the M2CVD (UniXcoder) configuration with Phase II integration surpasses its UniXcoder counterpart by 3.29%.
@@
Result Answer to RQ2: Experiments show that the vulnerability description refinement process of M2CVD can effectively improve the performance of code vulnerability detection.
5.3. RQ3. Effects of vulnerability description refinement for LLMs
@@
ChatGPT 45.29%
ChatGPT-fewshot 45.30%
@@
UniXcoder for CodeLLaMa-13B 52.59%
ALL“YES” for ChatGPT 50.04%
ALL ”NO” for ChatGPT 52.12%
UniXcoder for ChatGPT 57.61%
@@
Result Answer to RQ3: The strategy of informing LLMs with the insights from code models that are attuned to the specific dataset logic not only improves the performance but also highlights the potential of collaborative learning systems in code vulnerability detection. This process effectively enhances the accuracy of LLMs in the code vulnerability detection task within the framework of M2CVD, and also enhances the accuracy of LLMs in adding vulnerability semantics.
@@
b) UniXcoder: Also adopts the optimal configuration as documented in existing literature.
@@
The result from the experiment presented in Table 5 provides experimental result about the impact of combining different code models and LLMs within the M2CVD framework for predicting code vulnerabilities. The standalone models, CodeBERT and UniXcoder, establish a baseline with accuracy of 63.59% and 64.82%, respectively. The combination of CodeBERT with CodeLLaMa-13B results in a slight accuracy increase, reaching 64.38%. When UniXcoder is paired with CodeLLaMa-13B, there is a more noticeable improvement, with the accuracy climbing to 64.93%. These figures serve as a benchmark to assess the added value of integrating LLMs with code models. More substantial gains in accuracy are observed when ChatGPT is introduced to the mix. ChatGPT paired with CodeBERT yields an accuracy of 66.10%, while its combination with UniXcoder tops the table at 68.11%. This indicates that the ChatGPT model significantly enhances the performance of both code models, with the ChatGPT+UniXcoder configuration proving to be the most effective partnership in this experiment.
Result Answer to RQ4: The synergy between LLMs and code models in the M2CVD framework significantly enhances the precision of detecting code vulnerabilities. Specifically, the more superior performing LLMs and code models contributes to a more pronounced accuracy improvement in the M2CVD synergy mechanism.
@@
Model Accuracy
CodeBERT 63.59%
CodeLLaMa-13B+CodeBERT 64.38%
ChatGPT+CodeBERT 66.10%
UniXcoder 64.82%
CodeLLaMa-13B+UniXcoder 64.93%
ChatGPT+UniXcoder 68.11%
@@ 7. DISCUSSION
Version of ChatGPT. In its latest release, ChatGPT 4 offers much more powerful generation and understanding capabilities than ChatGPT 3.5. However, due to individual user usage limits, each person is currently only allowed to access the API 200 times. At the same time, it needs to pay a relatively expensive usage fee, which makes it difficult to generate the vulnerability semantics of tens of thousands of code fragments through ChatGPT 4. In comparison, ChatGPT 3.5 has a more relaxed access policy and price, but also a more powerful understanding and generation capability.
@@ 8. CONCLUSION
In this paper, we introduced M2CVD, a novel method designed to address the challenge of software vulnerability detection by harnessing the combined strengths of pre-trained code models and large language models. The M2CVD integrates the language models such as ChatGPT and code models like UniXcoder, to create a collaboration process capable of detecting vulnerabilities with high accuracy. Empirical evaluations conducted on the REVEAL and Devign datasets have demonstrated the effectiveness of M2CVD, showcasing its superior performance in detecting code vulnerabilities compared to existing benchmarks. The results of this research not only confirm the viability of M2CVD as a high-fidelity detection system but also underscore the potential of model synergy in enhancing the capabilities of automated vulnerability detection mechanisms. In essence, M2CVD demonstrates the potential to exploit the ability of different models to work together, providing a new idea for future research in automated software vulnerability detection and a scalable and effective solution for protecting software systems from changing threats.

### v2

@@ 5.2. RQ2: Effects of vulnerability description refinement for detection performance
b) UniXcoder-base: Also adopts the optimal configuration as documented in existing literature.
@@
d) M2CVD(UniXcoder-base) w/o PII: It omits the M2CVD comparison process. The LLM’ first vulnerability detection, when combined with the code, employs the UniXcoder-base model for prediction.
@@
f) M2CVD(UniXcoder-base): Also use the standard M2CVD process, with ChatGPT as the LLMs and UniXcoder-base as the code model.
g) M2CVD(GPT 4o + UniXcoder-nine): M2CVD experiments were performed on the latest ChatGPT 4o and code model UniXcoder-nine.
The results from the experiment as shown in the table 2, which presents a comparison results of model accuracy under different M2CVD configurations. The experimental result evidence suggests a improvement in model accuracy when integrating Phase II feedback into the M2CVD framework. Notably, the M2CVD (CodeBERT) w/o PII outperforms the CodeBERT model by a margin of 1.91%. Similarly, the M2CVD (UniXcoder-base) w/o PII configuration outperforms the UniXcoder by a margin of 2.23%. The enhancements are more pronounced when the comparison includes Phase II, as observed with the M2CVD (CodeBERT) configuration, which incorporates Phase II feedback, outperforms the CodeBERT model by a margin of 2.51%. And the M2CVD (UniXcoder-base) configuration with Phase II integration surpasses its UniXcoder-base counterpart by 3.29%.
The UniXcoder-nine model is obtained by continuing training on code-natural language pairs based on UniXcoder-base. GPT 4o provides an overly aggressive description of vulnerabilities without the refinement process, and considers almost all of the code to be vulnerable. This leads to the phenomenon that the model overfits on the training set and the vulnerability detection performance decreases.
@@
Answer to RQ2: Experiments show that the vulnerability description refinement process of M2CVD can effectively improve the performance of code vulnerability detection.
5.3. RQ3. Effects of hints of code models for LLMs
@@
ChatGPT 3.5 COT 45.29%
ChatGPT 3.5 3-fewshot 45.30%
ChatGPT 4o COT 53.73%
@@
UniXcoder-base for CodeLLaMa-13B 52.59%
ALL“YES” for ChatGPT 3.5 50.04%
ALL ”NO” for ChatGPT 3.5 52.12%
UniXcoder-base for ChatGPT 4o 56.73%
UniXcoder-base for ChatGPT 3.5 57.61%
@@
Answer to RQ3: The strategy of informing LLMs with the insights from code models that are attuned to the specific dataset logic not only improves the performance but also highlights the potential of collaborative learning systems in code vulnerability detection. This process effectively enhances the accuracy of LLMs in the code vulnerability detection task within the framework of M2CVD, and also enhances the accuracy of LLMs in adding vulnerability semantics.
@@
b) UniXcoder: Adopts the optimal configuration as documented in existing literature.
@@
The result from the experiment presented in Table 4 provides experimental result about the impact of combining different code models and LLMs within the M2CVD framework for predicting code vulnerabilities. The standalone models, CodeBERT and UniXcoder, establish a baseline with accuracy of 63.59% and 64.82%, respectively. The combination of CodeBERT with CodeLLaMa-13B results in a slight accuracy increase, reaching 64.38%. When UniXcoder is paired with CodeLLaMa-13B, there is a more noticeable improvement, with the accuracy climbing to 64.93%. These figures serve as a benchmark to assess the added value of integrating LLMs with code models. More substantial gains in accuracy are observed when ChatGPT is introduced to the mix. ChatGPT paired with CodeBERT yields an accuracy of 66.10%, while its combination with UniXcoder tops the table at 68.11%. The result from the experiment presented in fig 4 provides a more intuitive result about the impact of combining different code models and LLMs within the M2CVD framework for predicting code vulnerabilities.
This indicates that the ChatGPT model significantly enhances the performance of both code models, with the ChatGPT+UniXcoder configuration proving to be the most effective partnership in this experiment.
@@
Model Parameters Accuracy
CodeBERT 125M 63.59%
CodeLLaMa-13B+CodeBERT 13B+125M 64.38%
ChatGPT 3.5+CodeBERT -+125M 66.10%
UniXcoder-base 223M 64.82%
CodeLLaMa-13B+UniXcoder-base 13B+223M 64.93%
ChatGPT 3.5+UniXcoder-base -+223M 68.33%
ChatGPT 4o+UniXcoder-base -+223M 68.88%
ChatGPT 4o+UniXcoder-nine -+223M 69.11%
Answer to RQ4: The synergy between LLMs and code models in the M2CVD framework significantly enhances the precision of detecting code vulnerabilities. Specifically, the more superior performing LLMs and code models contributes to a more pronounced accuracy improvement in the M2CVD synergy mechanism.
@@ 7. Discussion
ChatGPT 4o Count ChatGPT 3.5-turbo Count
Memory Management Issues 21 Buffer Overflow 5
Input Validation Issues 18 Memory Leaks 5
Boundary and Overflow Issues 16 Improper Error Handling 5
Error Handling Issues 10 Integer Overflow 5
Concurrency and Synchronization Issues 4 Null Pointer Dereference 3
Table 5. Top 5 Vulnerability Types from ChatGPT 3.5 and ChatGPT 4o
Version of ChatGPT. In the latest release, ChatGPT 4o offers enhanced generation and understanding capabilities compared to ChatGPT 3.5. However, the relatively high usage fees associated with ChatGPT 4o make it impractical for generating vulnerability semantics for tens of thousands of code fragments. On the other hand, ChatGPT 3.5 has more lenient access policies and pricing, making it a more feasible option for large-scale tasks.
We performed M2CVD on the Devign dataset with different ChatGPT and UniXcoder-base. We conducted experiments using M2CVD on the Devign dataset, comparing different versions of ChatGPT and UniXcoder-base. The cost of using ChatGPT 4o for defect detection on the full dataset amounted to approximately \$1200. Despite the higher cost, the experimental results indicated that the vulnerability descriptions generated by ChatGPT 4o were not significantly better than those generated by ChatGPT 3.5. Consequently, ChatGPT 3.5 was selected as the default LLM version for M2CVD.
Our sampling analysis on the Devign dataset revealed that ChatGPT 4o reported 98% of its code as vulnerable due to its overly strict vulnerability definition, resulting in an F1 score of only 8.16%. After applying the M2CVD process, the F1 score improved to 19.23%.
Table 5 summarizes the vulnerability reporting by GPT 4o. The high false positive rate of the GPT 4o model can be attributed to several factors. The detector lacks global context information when analyzing code snippets, relying solely on the snippet itself for judgment. This limitation can lead to false positives, such as failing to recognize pre-initialized variables or pre-performed bounds checks in the calling function. The detector’s strict criteria flag even minor potential issues as vulnerabilities, like uninitialized variables or unchecked pointers, even when they are safe within the specific context. The detector is highly sensitive to boundary checking and input validation, resulting in numerous unnecessary warnings even in scenarios where out-of-bounds access is impossible. By assuming the worst-case scenario, the detector enhances code robustness in some instances but often imposes excessive error-handling logic that is not required in most real-world applications.
ChatGPT3.5’s vulnerability determination logic is not that aggressive. GPT 3.5 gives 70 defect judgments and 30 non-defect judgments for 100 sampled data. The accuracy is 53% and the F1 is 38.96%. After the process of M2CVD, ACC is increased to 55%, and F1 is greatly increased to 50.55%. GPT 3.5 gives 56 vulnerability judgments, which tend to follow the distribution of this data set. Considering the sampling experiment results, ChatGPT 3.5 is used as the default LLM in this method.
@@ 8. Conclusion
In this paper, we introduce M2CVD, a novel method designed to address the challenge of software vulnerability detection by harnessing the combined strengths of pre-trained code models and large language models. The M2CVD integrates the language models such as ChatGPT and code models like UniXcoder, to create a collaboration process capable of detecting vulnerabilities with high accuracy. Empirical evaluations conducted on the REVEAL and Devign datasets have demonstrated the effectiveness of M2CVD, showcasing its superior performance in detecting code vulnerabilities compared to existing benchmarks. The results of this research not only confirm the viability of M2CVD as a high-fidelity detection system but also underscore the potential of model synergy in enhancing the capabilities of automated vulnerability detection mechanisms. In essence, M2CVD demonstrates the potential to exploit the ability of different models to work together, providing a new idea for future research in automated software vulnerability detection and a scalable and effective solution for protecting software systems from changing threats.